A pool can be filled in 12 hours if water enters through a pipe alone or in 30 hours if water enters through a hose alone. If water is entering through the pipe and the hose, how long will it take to fill the pool?
Answer by richwmiller(17219)

x/12+x/30=1
5x+2x=60
7x=60
x=60/7=8.57 hrs